OpenStudy (anonymous):

Solve the system using elimination. 2x+6y=-12 5x-5y=10

OpenStudy (amorfide):

I will let @AbhimanyuPudi answer, i gotta go, if anything is wrong i shall correct when i return <3

OpenStudy (anonymous):

2x+6y=-12 2x=-12-6y x=-6-3y Substitute this value of x in the second equation, 5x-5y=10 5(-6-3y)-5y=10 -30-15y-5y=10 y=-2 Substituite y=-2 in any of the equation to find x..I'll substitute it in second equation 5x-5y=10 5x-5(-2)=10 5x+10=10 5x=10-10 5x=0 x=0 So, x=0 and y=-2

OpenStudy (anonymous):

Omg. Thanks soooo much!

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

wanna help with some more??(:

OpenStudy (anonymous):

sure

OpenStudy (amorfide):

that was the substitution method not elimination...

OpenStudy (anonymous):

ya....elimination is like this: multiply the first equation with 5 and the second equation with 2, you have 10x+30y= -60 and 10x-10y=20 Subtract the second one from the first, 10x+30y-10x+10y= -60-20 40y= -80 y= -2 Substitute this value in any of the given equations and you get the value of x x=0